OpenStudy (anonymous):

A card is drawn from a standard deck of cards. The card is not replaced, and a second one is drawn. Find the probaility. P(queen of hearts or a red card) P(queen of hearts or 10)

OpenStudy (kropot72):

Hypergeometric distribution applies

OpenStudy (anonymous):

i could be wrong on this

OpenStudy (anonymous):

oh i didnt read first part right

OpenStudy (anonymous):

26of 52 then 5 of 51 if didnt draw q hearts or red ten first time

OpenStudy (anonymous):

i dont get it

OpenStudy (kropot72):

Do you want to go thru the second part of the question?

OpenStudy (anonymous):

if u can :)

OpenStudy (kropot72):

There are 26 red cards including the queen of hearts and a total of 52 cards in a standard pack. The probability of getting a queen of hearts or a red card in a sample of two without replacement is given by: \[\frac{\left(\begin{matrix}26 \\ 1\end{matrix}\right)\left(\begin{matrix}26 \\ 1\end{matrix}\right)}{\left(\begin{matrix}52 \\ 2\end{matrix}\right)}\] \[=\frac{26}{51}\]

OpenStudy (kropot72):

There are 4 cards with value 10 and one queen of hearts card giving a total of 5. The probability of getting a queen of hearts or a 10 card in a sample of two without replacement is given by: \[\frac{\left(\begin{matrix}5 \\ 1\end{matrix}\right)\left(\begin{matrix}47 \\ 1\end{matrix}\right)}{\left(\begin{matrix}52 \\2\end{matrix}\right)}\] \[=\frac{235}{1326}\]